A random sample of 11 cigarettes of a certain brand has an average nicotine content of 6.6 milligrams and a standard deviation of 2.5 milligrams. Is this in line with the manufacturer's claim that the average nicotine content does not exceed 4.2 milligrams? Use 0.1 level of significance.


1
Expert's answer
2022-04-06T07:50:41-0400

"\mu=4.2, \ n=11, \ \bar{x}=6.6, \ s=2.5,\ \alpha=0.1."

The null and alternative hypotheses are

"H_0:\mu\le4.2,\\
H_1:\mu>4.2."

Because "\sigma" is unknown and he population is normally distributed, we use the t-test.

The test is a right-tailed test, the level of significance is "\alpha=0.1" , and the degrees of freedom are

d.f. = 11 - 1 = 10. So, using t-table, the critical value is t0 = 1.372. The rejection region is t > 1.372. The standardized test statistic is

"t=\cfrac{\bar{x}-\mu}{s/\sqrt{n}}=\cfrac{6.6-4.2}{2.5/\sqrt{11}}=3.184>1.372."

Because t is in the rejection region, we reject the null hypothesis, the average nicotine content exceeds 4.2 milligrams.